OpenStudy (anonymous):

Which of the following is a solution to (√x+10)+8=3? Answers: x = 15 x = 5 x = –5 x = –15

OpenStudy (angela210793):

sqrx+10=-5 x+10=25 x=15

OpenStudy (anonymous):

it was for the previous equation also

OpenStudy (akshay_budhkar):

root of x+10=root 25=5+8=13 so angela you are wrong.

OpenStudy (anonymous):

agree with Akshay

OpenStudy (anonymous):

it will be x = 225 cause the result is \[\sqrt{x}=-15\]

OpenStudy (akshay_budhkar):

root of anything cannot be negative

OpenStudy (akshay_budhkar):

the answer to it is no solution.. i can bet.

OpenStudy (anonymous):

right answer 225

OpenStudy (angela210793):

wht????????????????? subtract 8 both sides and u'll get sqr(x+10)=-5 raise to the 2nd power x+10=(-5)^2=25 subtract 10 x=25-10 x=15

OpenStudy (akshay_budhkar):

sqr(x+10)=-5 this is wrong. root cannot be negetive check the graph of root x

OpenStudy (angela210793):

sqr 25 may be -5 can't it?

OpenStudy (anonymous):

see the equation. not whole squroot only \[\sqrt{x}\]

OpenStudy (angela210793):

i know everything abt sqrx and its graph...

OpenStudy (akshay_budhkar):

no sqr 25 cannot be -5 check the graph

OpenStudy (akshay_budhkar):

its always mod

OpenStudy (angela210793):

well..why isn't thr any alternative no answer or smth like tht then? O.o

OpenStudy (akshay_budhkar):

no the answer is simply no solution if you go by rules if the chapter or topic demands what you have done then it is right mathematically its wrong

OpenStudy (angela210793):

i get ur point :)